Find the area of region y =17x2/16, y = 9x2/4 and y = 1 - 4x, satisfying x ≤0.


1
Expert's answer
2022-05-03T10:13:08-0400
"\dfrac{17x^2}{16}=1-4x, x\le0"

"17x^2+64x-16=0"

"(17x-4)(x+4)=0"

Since "x\le0," we take "x=-4."


"\dfrac{9x^2}{4}=1-4x, x\le0"

"9x^2+16x-4=0"

"(9x-2)(x+2)=0"

Since "x\le0," we take "x=-2."


"\dfrac{17x^2}{16}=\dfrac{9x^2}{4}"

"x=0"

"A=\displaystyle\int_{-4}^{-2}(1-4x-\dfrac{17x^2}{16})dx"

"+\displaystyle\int_{-2}^{0}(\dfrac{9x^2}{4}-\dfrac{17x^2}{16})dx"

"=[x-2x^2-\dfrac{17x^3}{48}]\begin{matrix}
 -2\\
 -4
\end{matrix}+[\dfrac{19x^3}{48}]\begin{matrix}
 0\\
 -2
\end{matrix}"

"=-2-8+\dfrac{17}{6}+4+32-\dfrac{68}{3}+\dfrac{19}{6}"

"=\dfrac{28}{3}({units}^2)"

Area= "\dfrac{28}{3}" square units.